# Sculpting Molecules in 3D: A Flexible Substructure Aware Framework for Text-Oriented Molecular Optimization

### v1

@@ Abstract
The integration of deep learning, particularly AI-Generated Content, with high-quality data derived from ab initio calculations has emerged as a promising avenue for transforming the landscape of scientific research. However, the challenge of designing molecular drugs or materials that incorporate multi-modality prior knowledge remains a critical and complex undertaking. Specifically, achieving a practical molecular design necessitates not only meeting the diversity requirements but also addressing structural and textural constraints with various symmetries outlined by domain experts. In this article, we present an innovative approach to tackle this inverse design problem by formulating it as a multi-modality guidance generation/optimization task. Our proposed solution involves a textural-structure alignment symmetric diffusion framework for the implementation of molecular generation/optimization tasks, namely 3DToMolo. 3DToMolo aims to harmonize diverse modalities, aligning them seamlessly to produce molecular structures adhere to specified symmetric structural and textural constraints by experts in the field. Experimental trials across three guidance generation settings have shown a superior hit generation performance compared to state-of-the-art methodologies. Moreover, 3DToMolo demonstrates the capability to generate novel molecules, incorporating specified target substructures, without the need for prior knowledge. This work not only holds general significance for the advancement of deep learning methodologies but also paves the way for a transformative shift in molecular design strategies. 3DToMolo creates opportunities for a more nuanced and effective exploration of the vast chemical space, opening new frontiers in the development of molecular entities with tailored properties and functionalities.
@@ 1 Introduction
In general, it is highly desirable to develop a methodology that is purposefully tailored for optimizing both 2D (atom types and chemical bond topology) and conformer structure (3D) aspects of molecules. Simultaneously, this methodology should exhibit compatibility with a broad spectrum of complex goals, facilitating multi-goal guidance optimization. Capitalizing on the remarkable capabilities exhibited by large language models (LLMs), there is a natural inclination to explore the feasibility of consolidating property and structure descriptions into a unified text format. Then, we are able to leverage the prowess of LLMs to extract a unified representation from such textual amalgamation. In pursuit of this, we advocate the training of a joint molecule diffusion model designed to capture the fine-grained distributions of 2D+3D molecule structures. The crux of an ideal molecule optimization lies in achieving alignment within the representation spaces of both the text side and the molecule structure side. To this end, we introduce the 3D-based Text-oriented Molecular Optimization (3DToMolo), wherein this specific cross-modality alignment is realized through contrastive training. This involves training the representation pair obtained from a lightweight LLM and an (SE(3)) equivariant graph transformer specifically tailored for molecules. The intermediary steps introduced during the forward diffusion process play a crucial role as a medium connecting the initial molecules with those possessing target properties. In contrast to generative approaches of sampling from white noise, the intermediate molecule representations retain essential structural information from the original molecules. Moreover, control over text descriptions is meticulously exerted at each step during the subsequent backward optimization process. Beyond the flexibility inherent in optimizing entire regions of molecules, our framework showcases its prowess in two practical scenarios where substructures are preserved. In these instances, specific three-dimensional structures are pre-fixed, and optimization exclusively occurs within the remaining inpainting areas, highlighting the versatility and effectiveness of our approach.
@@ 2.1 Definition of text - structural optimization
Figure 1: Overview of 3DToMolo. a. The alignment of textual description and chemical structures of molecules, which is realized through contrastive learning of the two latent representations: molecule structure encoding with its paired text embedding. b. Conditional diffusion model. In order to maintain molecule optimization in alignment with the prompt, conditional diffusion model incorporates text prompts at each step during the subsequent backward optimization process. c. The zero-shot prompt-driven molecule optimization task involves modifying the input molecule in response to a given text prompt related to physicochemical properties. 3DToMolo necessitates the overall optimization of both 2D and 3D features of molecules, ensuring a balanced alignment with the input molecule and the text prompt which is achieved by the conditional diffusion model, shown as b. d. Molecule optimization under structural constraints. This task further enhances the similarity to the input molecule by retaining essential structural features. e. Molecule optimization under appointed sites. Given the precise position within the input molecule, 3DToMolo aims to optimize molecule by offering strategies for atoms and the bonds connected with the site.
@@ 2.2 Development of a text-structural diffusion model
On the prompt-text embedding side, we leverage the widely acclaimed large language model, LLAMA [37], as the text encoder, tapping into its ability to capture nuanced semantic representations from textual descriptions. Then, the alignment is achieved through contrastive learning of the two latent representations: the molecule structure encoding and its paired text embedding. As a possible extension, the text-structure alignment can be independently fine-tuned for domain-specific texts, e,g., materials . To validate the effectiveness of our learned molecule latent embedding, we conduct tests on retrieval and property prediction tasks. The experimental results are provided in Table 1. In line with prior research on molecule pretraining [38, 11], we adopt the MoleculeNet benchmark [39], which encompasses eight single-modal binary classification datasets aimed at evaluating the efficacy of pretrained molecule representation approaches. We adopt the area under the receiver operating characteristic curve (ROC-AUC) [40] as the evaluation metric. As delineated in Table 1, our observations reveal that methods based on pretraining markedly enhance overall classification accuracy compared to randomly initialized counterparts. Additionally, 3DToMolo demonstrates superior performance on five out of eight tasks, while achieving comparable results to the leading baselines in the remaining three tasks. Since 3DToMolo lies in its ability to leverage pretrained chemical structure representations that incorporate external domain knowledge, which potentially provides a beneficial implicit bias for property prediction tasks. The key hyperparameters of molecule encoder are layers of Graph Transformer $5$, learning rate $\{1e-4,1e-5\}$, hidden states dimension (X: 256, E: 128, pos: 64).
@@
By integrating the alignment objective into the denoising process (with the noising step $t$ iteratively tuned to retain the similarity ratio between $M_{y}$ and $M_{0}$), as detailed in the method section, the model is capable of optimizing molecules with desired properties in a seamless, end-to-end manner. Notably, 3DToMolo is zero-shot, signifying that throughout the optimization process. We refrain from introducing any feedback for multi-stage correction, and the denoising process is executed in a single run. We propose three types of downstream tasks in the following sections and systematically verify the robust effectiveness of 3DToMolo for text-structural optimization.
@@ 2.3 Flexible molecule optimization under Physicochemical Property Prompts
  • GPT3.5: With its immense language processing capabilities and a broad understanding of chemistry concepts, has the potential to revolutionize molecule optimization.
  • Galactic: A versatile scientific language model, extensively trained on a vast repository of scientific text and data.
@@
Table 2: The results of Molecule optimization on the PubChemSTM dataset. Best baseline results are highlighted with underlined text. Best overall results are marked by *. Statistically significant improvement (t-test over 5 different dataset splits, p-value$<$ 0.05) is highlighted with bold text.
Prompts
Baselines
Ours
MoleculeSTM
GPT3.5
Galactic
This molecule has low HOMO (Highest occupied molecular orbital) value, which is more stable.
35.00
37.50
00.00
$\textbf{46.50}^{*}$
This molecule has high HOMO (Highest occupied molecular orbital) value, which is more reactive and susceptible to electron acceptance or participation in chemical reactions.
28.00
40.50
00.00
$\textbf{58.50}^{*}$
This molecule has low LUMO (Lowest Unoccupied Molecular Orbital) value.
09.50
27.50
00.00
$\textbf{30.00}^{*}$
This molecule has high LUMO (Lowest Unoccupied Molecular Orbital) value.
53.50
30.50
00.00
$\textbf{88.00}^{*}$
This molecule has low HOMO-LUMO gap value, which has enhanced light absorption properties. The small energy difference allows the molecule to absorb photons in the visible or ultraviolet range, resulting in a higher likelihood of exhibiting color or being used as a dye or pigment.
09.00
29.50
00.00
$\textbf{87.50}^{*}$
This molecule has high HOMO-LUMO gap value, which is insulating or non-conductive. The large energy difference between the HOMO and LUMO orbitals makes it less likely for electrons to be excited across the gap, resulting in low electrical conductivity.
56.50
40.00
00.00
$\textbf{89.00}^{*}$
This molecule has high polarity.
41.50
44.50
00.00
$\textbf{58.00}^{*}$
This molecule has low polarity.
45.00
35.50
00.00
$\textbf{73.50}^{*}$
This molecule is soluble in water, which may have high polarity.
28.50
25.00
04.00
$\textbf{46.00}^{*}$
This molecule is insoluble in water, which may have low polarity.
52.00
74.00
03.50
$\textbf{81.00}^{*}$
This molecule is soluble in water.
$\underline{29.50}^{*}$
24.50
04.00
26.50
This molecule is insoluble in water.
52.00
66.50
03.50
$\textbf{88.00}^{*}$
This molecule has high permeability.
34.50
23.00
05.50
$\textbf{89.00}^{*}$
This molecule has low permeability.
24.50
$\textbf{39.00}^{*}$
00.50
21.00
This molecule has more hydrogen bond acceptors.
07.50
$\underline{20.00}^{*}$
00.00
10.50
This molecule has more hydrogen bond donors.
05.00
12.50
00.00
$\textbf{34.50}^{*}$
This molecule is like a drug.
42.00
$\underline{59.00}^{*}$
00.00
41.50
This molecule is not like a drug.
39.00
31.00
00.00
$\textbf{56.50}^{*}$
@@
Visual analysis on multi-objective molecule optimizations. We further analyze the multi-objective molecule optimization. Water solubility and polarity are two positively correlated properties. Consequently, 3DToMolo turns the 2-oxo-1-pyrindinyl group into a benzene group, which reduces the solubility as well as the polarity of the input molecule (Figure 2(f1)). In contrast, 3DToMolo adds a hydroxyl group to the input molecule when given the opposite prompt, which increases the solubility and the polarity (Figure 2(f2)). More results on multi-objective optimization are presented in Supplementary Table S5. We observe that in the multi-objective task of improving both the solubility and the polarity, $46\%$ (92 out of 200) of the input molecules have been observed a solubility improvement. This ratio is higher than the hit ratio of the single-objective solubility improvement task, which is $26.5\%$ as shown in Table 2. It hints that coupling with the polarity in the prompt helps us better tune the solubility. A possible reason could be that 3DToMolo tunes the polarity more flexibly, as discussed in the next paragraph.
@@
Figure 5: Molecule optimization with structural constraints. (a) Unconditional modification on the input molecular skeleton. The red-orange regions mark the preserved skeleton in output molecules. (b) A case study for optimization on the internal region. On the left is a reported spiro-linked $\pi-$conjugated molecule, tetraphenylsilane, used as the template. The center silicon atom (yellow-shaded area) is diffused while the four benzene rings remain fixed until final steps of the denoising process. On the right are two selected output structures whose benzene rings remain non-coplanar as desired. Hydrogen atoms are not shown for the sake of clear visualization. (c) Redox-related prompt-driven molecule optimization with preserved skeletons.
2.4 Molecule optimization with Structural constraints
@@
In the following case studies, we have carefully chosen specific representative molecules to conduct a thorough examination. Specifically, we designate the substructure to be protected, denoted as $S_{0}$, to encompass all core structures excluding the hydrogen atoms or other removable atoms. By only optimizing the pre-defined removable atoms, we aim to maintain the integrity of the molecular backbone, emphasizing the impact on non-hydrogen constituents, which play key roles in defining the original molecule’s chemical properties and functionality.
Unconditional molecule optimization with structure constraints. We select three molecules for unconditional optimization while preserving their skeletons. As seen in Figure 5(a), the output molecules exhibit a great variety in modifications, including generation of diverse heteroatoms, simultaneous substitutions on multi-sites, formation of cyclic structures, extension of aromatic rings and so on. Meanwhile, the seed skeletons are preserved successfully. We note that some of the complex outputs can hardly be synthesized from the input molecule. However, such unconditional optimization can be utilized as a tool to extensively search the chemical space for molecules that have the matching skeleton. The structure of the desired skeleton can be passed on to the model in the appearance of input molecule. In this sense, we should not be bothered much by the synthetic path from the input molecule to the optimization results.
Internal-region molecule optimization. Modification on the internal region of a molecule is challenging [44], as it necessitates rational linkages of fragments. This task becomes even more difficult when specific stereochemistry requirements are imposed. Here, we demonstrate the competence of 3DToMolo for such tasks via a case study on tetraphenylsilane. Non-coplanar benzene rings in tetraphenylsilane (Figure 5(b), left) is a desired structural feature for optical materials that have high refractive index and low light double-refraction (birefringence) [45, 46]. Benzene rings contribute to the strong refractive ability. The non-coplanar configuration hinders the $\pi-\pi$ stacking, preventing the formation of layering structure and thus reducing the double-refraction. To generate more candidate molecules with satisfactory configuration, we spatially fix the benzene rings in the tetraphenylsilane molecule and diffuse the center silicon atom. The structural constraint is lifted at final steps during the denoising process. Valid generated structures are examined by Density Functional Theory (DFT) [47, 48] computation. Most of the structures have connected four benzene rings via the generated central motif and maintain the rings non-coplanar. Two optimized results are exemplified in Figure 5(b) and more can be found in Supplementary Figure S3. As a comparison, GPT3.5’s performance on the same task is poor either because it fails to generate required structures or because it merely conducts single-atom replacement on the silicon atom without the capability of providing more sophisticated internal structures (Supplementary Section S8).
@@
Redox potential related molecule optimization. 3DToMolo extends its applicability to broader physicochemical tasks, exemplified here by its performance in optimizing the redox potential of input molecules. In the context of energy storage, enhancing the energy density of batteries necessitates elevated voltage, requiring electrolyte molecules with an expansive electrochemical window. As case studies, we aim to increase the oxidation potential and decrease the reduction potential of exemplary electrolyte molecules.
@@
Thiophene is a common structure used in electrolyte additives for lithium-ion batteries. To augment thiophene’s resistance to high voltage, we apply the prompt "This molecule has low HOMO (Highest occupied molecular orbital) value, which has high oxidation potential" while constraining all atoms except hydrogens. A comparative experiment without the prompt serves as the baseline. With the prompt, 19.5% of the generated derivatives exhibit an increased oxidation potential, compared to 12.3% without the prompt. Select successful examples are presented in Figure 5(b). A similar experiment is conducted on phosphate, frequently employed in electrolytes to enhance battery stability at elevated temperatures. The success rate is 8.12% with the prompt and 5.66% without. To illustrate the modification of reduction potential, we use quinoxaline, pertinent to redox flow batteries. By constraining all atoms except hydrogens and employing corresponding prompts, we successfully modify the reduction potential in two directions (Figure 5(b)). In the more desirable direction of lowering the reduction potential, we achieve a success rate of 69.1%.

### v2

# Sculpting Molecules in Text-3D Space: A Flexible Substructure Aware Framework for Text-Oriented Molecular Optimization

@@ Abstract
The integration of deep learning, particularly AI-Generated Content, with high-quality data derived from ab initio calculations has emerged as a promising avenue for transforming the landscape of scientific research. However, the challenge of designing molecular drugs or materials that incorporate multi-modality prior knowledge remains a critical and complex undertaking. Specifically, achieving a practical molecular design necessitates not only meeting the diversity requirements but also addressing structural and textural constraints with various symmetries outlined by domain experts. In this article, we present an innovative approach to tackle this inverse design problem by formulating it as a multi-modality guidance optimization task. Our proposed solution involves a textural-structure alignment symmetric diffusion framework for the implementation of molecular optimization tasks, namely 3DToMolo. 3DToMolo aims to harmonize diverse modalities including textual description features and graph structural features, aligning them seamlessly to produce molecular structures adhere to specified symmetric structural and textural constraints by experts in the field. Experimental trials across three guidance optimization settings have shown a superior hit optimization performance compared to state-of-the-art methodologies. Moreover, 3DToMolo demonstrates the capability to discover potential novel molecules, incorporating specified target substructures, without the need for prior knowledge. This work not only holds general significance for the advancement of deep learning methodologies but also paves the way for a transformative shift in molecular design strategies. 3DToMolo creates opportunities for a more nuanced and effective exploration of the vast chemical space, opening new frontiers in the development of molecular entities with tailored properties and functionalities.
@@ 1 Introduction
In general, it is highly desirable to develop a methodology that is purposefully tailored for optimizing both 2D (atom types and chemical bond topology) and conformer structure (3D) aspects of molecules. Simultaneously, this methodology should exhibit compatibility with a broad spectrum of complex goals, facilitating multi-goal guidance optimization. Capitalizing on the remarkable capabilities exhibited by large language models (LLMs), there is a natural inclination to explore the feasibility of consolidating property and structure descriptions into a unified text format. Then, we are able to leverage the prowess of LLMs to extract a unified representation from such textual amalgamation. In pursuit of this, we advocate the training of a joint molecule diffusion model designed to capture the fine-grained distributions of 2D+3D molecule structures. The crux of an ideal molecule optimization lies in achieving alignment within the representation spaces of both the text side and the molecule structure side. To this end, we introduce the 3D-based Text-oriented Molecular Optimization (3DToMolo), wherein this specific cross-modality alignment is realized through contrastive training. This involves training the representation pair obtained from a lightweight LLM and an (SE(3)) equivariant graph transformer specifically tailored for molecules. The intermediary steps introduced during the forward diffusion process play a crucial role as a medium connecting the initial molecules with those possessing target properties. In contrast to generative approaches of sampling from white noise, the intermediate molecule representations retain essential structural information from the original molecules. Moreover, control over text descriptions is meticulously exerted at each step during the subsequent backward optimization process. Beyond the flexibility inherent in optimizing entire regions of molecules, 3DToMolo showcases its prowess in two practical scenarios where substructures are preserved. In these instances, specific three-dimensional structures are pre-fixed, and optimization exclusively occurs within the remaining inpainting areas, highlighting the versatility and effectiveness of 3DToMolo.
@@ 2.1 Definition of text - structural optimization
Figure 1: Overview of 3DToMolo. (a). The alignment of textual description and chemical structures of molecules, which is realized through contrastive learning of the two latent representations: molecule structure encoding with its paired text embedding. (b). Conditional diffusion model. In order to maintain molecule optimization in alignment with the prompt, conditional diffusion model incorporates text prompts at each step during the subsequent backward optimization process. (c). The zero-shot prompt-driven molecule optimization task involves modifying the input molecule in response to a given text prompt related to physicochemical properties. 3DToMolo necessitates the overall optimization of both 2D and 3D features of molecules, ensuring a balanced alignment with the input molecule and the text prompt which is achieved by the conditional diffusion model, shown as (b). (d). Molecule optimization under structural constraints. This task further enhances the similarity to the input molecule by retaining essential structural features. (e). Molecule optimization under appointed sites. Given the precise position within the input molecule, 3DToMolo aims to optimize molecule by offering strategies for atoms and the bonds connected with the site.
@@ 2.2 Development of a text-structural diffusion model
On the prompt-text embedding side, we leverage the widely acclaimed large language model, LLAMA [37], as the text encoder, tapping into its ability to capture nuanced semantic representations from textual descriptions. Then, the alignment is achieved through contrastive learning of the two latent representations: the molecule structure encoding and its paired text embedding. As a possible extension, the text-structure alignment can be independently fine-tuned for domain-specific texts, e,g., materials [38]. To validate the effectiveness of our learned molecule latent embedding, we conduct tests on retrieval and property prediction tasks. The experimental results are provided in Table 1. In line with prior research on molecule pretraining [39, 11], we adopt the MoleculeNet benchmark [40], which encompasses eight single-modal binary classification datasets aimed at evaluating the efficacy of pretrained molecule representation approaches. We consider nine pretraining-based methods as baselines: AttrMask [39], ContextPred [39], InfoGraph [41], MolCLR [42], GraphMVP [10], MoleculeSTM [11], GEM [43], and MGSSL (BFS/DFS) [44]. We adopt the area under the receiver operating characteristic curve (ROC-AUC) [45] as the evaluation metric. As delineated in Table 1, our observations reveal that methods based on pretraining markedly enhance overall classification accuracy compared to randomly initialized counterparts. Additionally, 3DToMolo demonstrates superior performance on five out of eight tasks, while achieving comparable results to the leading baselines in the remaining three tasks. Since 3DToMolo lies in its ability to leverage pretrained chemical structure representations that incorporate external domain knowledge, which potentially provides a beneficial implicit bias for property prediction tasks. The key hyperparameters of molecule encoder are layers of Graph Transformer $5$, learning rate $\{1e-4,1e-5\}$, hidden states dimension (X: 256, E: 128, pos: 64).
@@
By integrating the alignment objective into the denoising process (with the noising step $t$ iteratively tuned to retain the similarity ratio between $M_{y}$ and $M_{0}$), as detailed in the method section, the model is capable of optimizing molecules with desired properties in a seamless, end-to-end manner. Notably, 3DToMolo is zero-shot, signifying that throughout the optimization process. We refrain from introducing any feedback for multi-stage correction, and the denoising process is executed in a single run. We propose three types of downstream tasks in the following sections and systematically verify the robust effectiveness of 3DToMolo for text-structural optimization. All downstream tasks conducted are based on zero-shot optimization, which refers to a scenario where a model is required to optimize or adapt to new tasks or data without having been explicitly trained on those specific tasks or data beforehand. For instance, in molecular optimization, a model might be able to modify a molecule to achieve a desired biological activity, even if it has never been trained on molecules with that specific task, by leveraging its understanding of chemical space learned during pretraining. Zero-shot optimization is particularly valuable because it allows for the exploration of new chemical spaces or the design of molecules with novel properties, without requiring a model to be retrained or fine-tuned on every new task.
@@ 2.3 Flexible molecule optimization under Physicochemical Property Prompts
  • GPT-3.5: With its immense language processing capabilities and a broad understanding of chemistry concepts, has the potential to revolutionize molecule optimization.
  • Galactica: A versatile scientific language model, extensively trained on a vast repository of scientific text and data.
@@
Table 2: Results on 18 text prompts oriented towards diverse physicochemical objectives. The inputs consist of 200 molecules randomly sampled from ZINC, with the evaluation measured by the hit ratio (%) of the property changes in each experiment. Best baseline results are highlighted with underlined text. Best overall results are marked by *. Statistically significant improvement (t-test over 5 different dataset splits, p-value$<$ 0.05) is highlighted with bold text.
Prompts Baselines 3DToMolo
MoleculeSTM GPT-3.5 Galactica
This molecule has low HOMO (Highest occupied molecular orbital) value, which is more stable. 35.00 37.50 00.00 $\textbf{46.50}^{*}$
This molecule has high HOMO (Highest occupied molecular orbital) value, which is more reactive and susceptible to electron acceptance or participation in chemical reactions. 28.00 40.50 00.00 $\textbf{58.50}^{*}$
This molecule has low LUMO (Lowest Unoccupied Molecular Orbital) value. 09.50 27.50 00.00 $\textbf{30.00}^{*}$
This molecule has high LUMO (Lowest Unoccupied Molecular Orbital) value. 53.50 30.50 00.00 $\textbf{88.00}^{*}$
This molecule has low HOMO-LUMO gap value, which has enhanced light absorption properties. The small energy difference allows the molecule to absorb photons in the visible or ultraviolet range, resulting in a higher likelihood of exhibiting color or being used as a dye or pigment. 09.00 29.50 00.00 $\textbf{87.50}^{*}$
This molecule has high HOMO-LUMO gap value, which is insulating or non-conductive. The large energy difference between the HOMO and LUMO orbitals makes it less likely for electrons to be excited across the gap, resulting in low electrical conductivity. 56.50 40.00 00.00 $\textbf{89.00}^{*}$
This molecule has high polarity. 41.50 44.50 00.00 $\textbf{58.00}^{*}$
This molecule has low polarity. 45.00 35.50 00.00 $\textbf{73.50}^{*}$
This molecule is soluble in water, which may have high polarity. 28.50 25.00 04.00 $\textbf{46.00}^{*}$
This molecule is insoluble in water, which may have low polarity. 52.00 74.00 03.50 $\textbf{81.00}^{*}$
This molecule is soluble in water. $\underline{29.50}^{*}$ 24.50 04.00 26.50
This molecule is insoluble in water. 52.00 66.50 03.50 $\textbf{88.00}^{*}$
This molecule has high permeability. 34.50 23.00 05.50 $\textbf{89.00}^{*}$
This molecule has low permeability. 24.50 $\underline{39.00}^{*}$ 00.50 21.00
This molecule has more hydrogen bond acceptors. 07.50 $\underline{20.00}^{*}$ 00.00 10.50
This molecule has more hydrogen bond donors. 05.00 12.50 00.00 $\textbf{34.50}^{*}$
This molecule is like a drug. 42.00 $\underline{59.00}^{*}$ 00.00 41.50
This molecule is not like a drug. 39.00 31.00 00.00 $\textbf{56.50}^{*}$
@@
Visual analysis on multi-objective molecule optimizations. We further analyze the multi-objective molecule optimization. Water solubility and polarity are two positively correlated properties. Consequently, 3DToMolo turns the 2-oxo-1-pyrindinyl group into a benzene group, which reduces the solubility as well as the polarity of the input molecule (Figure 2(f1)). In contrast, 3DToMolo adds a hydroxyl group to the input molecule when given the opposite prompt, which increases the solubility and the polarity (Figure 2(f2)). More results on multi-objective optimization are presented in Supplementary Table S5. We observe that in the multi-objective task of improving both the solubility and the polarity, $46\%$ of the input molecules have been observed a solubility improvement after optimization, while $31.5\%$ have improvements in both properties, higher than the hit ratio of the single-objective solubility improvement task (Table 2). It hints that coupling with the polarity in the prompt helps us better tune the solubility. A possible reason could be that 3DToMolo tunes the polarity more flexibly, as discussed in the next paragraph.
@@
Figure 5: Molecule optimization with structural constraints. (a) Redox-related prompt-driven molecule optimization with preserved skeletons. (b) A case study for optimization on the internal region. On the left is a reported spiro-linked $\pi-$conjugated molecule, tetraphenylsilane, used as the template. The center silicon atom (yellow-shaded area) is diffused while the four benzene rings remain fixed until final steps of the denoising process. On the right are two selected output structures whose benzene rings remain non-coplanar as desired. Hydrogen atoms are not shown for the sake of clear visualization.
2.4 Prompt-driven Molecule optimization with Structural constraints
@@
In the following experiments, we have carefully chosen specific physicochemical tasks to conduct a thorough examination. Specifically, we protect all core structures excluding the hydrogen atoms or other removable atoms. By only optimizing the pre-defined removable atoms, we aim to maintain the integrity of the molecular backbone, emphasizing the impact on non-hydrogen constituents, which play key roles in defining the original molecule’s chemical properties and functionality.
@@
Redox potential related prompt-driven molecule optimization. In the context of energy storage, enhancing the energy density of batteries necessitates elevated voltage, requiring electrolyte molecules with an expansive electrochemical window. As case studies, we aim to increase the oxidation potential and decrease the reduction potential of exemplary electrolyte molecules.
@@
Thiophene is a common structure used in electrolyte additives for lithium-ion batteries. To augment thiophene’s resistance to high voltage, we apply the prompt "This molecule has low HOMO (Highest occupied molecular orbital) value, which has high oxidation potential" while constraining all atoms except hydrogens. A comparative experiment without the prompt serves as the baseline. With the prompt, 19.5% of the generated derivatives exhibit an increased oxidation potential, compared to 12.3% without the prompt. Selected successful examples are presented in Figure 5(a). A similar experiment is conducted on phosphate, frequently employed in electrolytes to enhance battery stability at elevated temperatures. The success rate is 8.12% with the prompt and 5.66% without. To illustrate the modification of reduction potential, we use quinoxaline, pertinent to redox flow batteries. By constraining all atoms except hydrogens and employing corresponding prompts, we successfully modify the reduction potential in two directions (Figure 5(a)). In the more desirable direction of lowering the reduction potential, we achieve a success rate of 69.1%.
Additionally, we conducted quantitative experiments a dateset of 143 common electrolyte additives [49]. We focus on properties people generally concern about in liquid electrolytes, including the redox potential, the polarity, and the water solubility. As shown in Table 3, 3DToMolo significantly outperforms the state-of-the-art baseline GPT-3.5. The major challenge for GPT-3.5 lies in the validity of output molecules, which contains problems such as incomplete rings and chemically incorrect bondings. 3DToMolo circumvents this defect, as the incorporation of 3D information makes it less likely to forget those hanging rings or bonding condition of atoms.
Table 3: The results of molecule optimization with structural constraints on the electrolyte additive dataset [49] were evaluated using the hit ratio (%) of property changes in each objective-oriented prompt. Statistically significant improvement (t-test over 5 different dataset splits, p-value$<$ 0.05) is highlighted with bold text.
Prompts Validity (%) Hit Ratio (%)
GPT-3.5 3DToMolo GPT-3.5 3DToMolo
This molecule has low HOMO (Highest occupied molecular orbital) value, which has high oxidation potential. 2.867 41.54 0.2797 10.91
This molecule has high HOMO (Highest occupied molecular orbital) value, which has low oxidation potential. 0.4196 42.73 0.06993 29.65
This molecule has low LUMO (Lowest Unoccupied Molecular Orbital) value, which has high reduction potential. 0 44.55 0 25.38
This molecule has high LUMO (Lowest Unoccupied Molecular Orbital) value, which has low reduction potential. 0.4895 42.45 0.4895 18.81
This molecule has high polarity. 1.399 42.03 1.119 30.35
This molecule has low polarity. 0.1399 43.15 00.00 12.17
This molecule is soluble in water. 2.448 55.24 1.678 36.50
This molecule is insoluble in water. 1.399 57.27 1.119 20.63
  • Note that the prompts we input into GPT-3.5 have subtle differences from those listed in the above table. To ensure a fair comparison, we formulated our request to GPT-3.5 as follows: "Optimize this molecule: <SMILES> to achieve <specified properties>. Optimization is restricted to hydrogen atoms only, while the rest of the input molecule remains fixed. The number of atoms (including hydrogen) added to the optimized molecule must be within the range of 1 to 10. Provide 10 different and valid results." For the compound prompts, the hit ratio is calculated based on the satisfaction of the redox potential part. In most cases, GPT-3.5 fails to meet the requirements, as the generated molecules do not conform to either the specified structural criteria or the desired properties.
Internal-region molecule optimization. Modification on the internal region of a molecule is challenging [50], as it necessitates rational linkages of fragments. This task becomes even more difficult when specific stereochemistry requirements are imposed. Here, we demonstrate the competence of 3DToMolo for such tasks via a case study on tetraphenylsilane. Non-coplanar benzene rings in tetraphenylsilane (Figure 5(b), left) is a desired structural feature for optical materials that have high refractive index and low light double-refraction (birefringence) [51, 52]. Benzene rings contribute to the strong refractive ability. The non-coplanar configuration hinders the $\pi-\pi$ stacking, preventing the formation of layering structure and thus reducing the double-refraction. To generate more candidate molecules with satisfactory configuration, we protect the benzene rings in the tetraphenylsilane molecule and diffuse the center silicon atom under the prompt "This molecule has four non-planar benzene rings". The protection is removed at final steps during the denoising process. Valid generated structures are examined by Density Functional Theory (DFT) [53, 54] computation. Most of the structures have connected four benzene rings via the generated central motif and maintain the rings non-coplanar. Two optimized results are exemplified in Figure 5(b) and more can be found in Supplementary Figure S4. As a comparison, GPT-3.5’s performance on the same task is poor either because it fails to generate required structures or because it merely conducts single-atom replacement on the silicon atom without the capability of providing more sophisticated internal structures (Supplementary Section S9).
We focus on the prompt-driven experiments in this section. However, we note that 3DToMolo is able to handle prompt-free unconditional generation (examples in Supplementary Section S8 and Figure S5), no different from other generative models that do not incorporate LLM models.